# Tuning Vision-Language Models with Candidate Labels by Prompt Alignment

### v2

@@ Abstract
Vision-language models (VLMs) can learn high-quality representations from a large-scale training dataset of image-text pairs. Prompt learning is a popular approach to fine-tuning VLM to adapt them to downstream tasks. Despite the satisfying performance, a major limitation of prompt learning is the demand for labelled data. In real-world scenarios, we may only obtain candidate labels (where the true label is included) instead of the true labels due to data privacy or sensitivity issues. In this paper, we provide the first study on prompt learning with candidate labels for VLMs. We empirically demonstrate that prompt learning is more advantageous than other fine-tuning methods, for handling candidate labels. Nonetheless, its performance drops when the label ambiguity increases. In order to improve its robustness, we propose a simple yet effective framework that better leverages the prior knowledge of VLMs to guide the learning process with candidate labels. Specifically, our framework disambiguates candidate labels by aligning the model output with the mixed class posterior jointly predicted by both the learnable and the handcrafted prompt. Besides, our framework can be equipped with various off-the-shelf training objectives for learning with candidate labels to further improve their performance. Extensive experiments demonstrate the effectiveness of our proposed framework.
@@
Vision-Language Models Prompt Learning Candidate Labels
@@ 1 Introduction
Large-scale vision-language models (VLMs) such as CLIP [38], ALIGN [19], and Coca [51] have become excellent base models in multiple domains, most of which employ a dual-encoder architecture to align the natural images with descriptive texts. Remarkably, this special training manner has endowed VLMs with superior zero-shot transfer performance on visual recognition tasks. In specific, during the inference, the pre-trained text encoder receives inputs in the form of man-crafted prompts, e.g., “a photo of <CLS>.”. Subsequently, all the generated textual embeddings are matched with the visual embedding obtained from the image encoder to predict the image category. However, the powerful zero-shot ability of VLMs was shown to be heavily dependent on the choice of handcrafted prompts, which needs substantial efforts and professional domain knowledge to design [57]. To avoid the manual design of the prompts, prompt learning [57] is proposed, which treats the textual prompt as additional learnable parameters and tunes them while keeping all the parameters of the pre-trained model fixed. Later, the concept of the prompt is extended to visual prompt [20] and multi-modal prompt [22] in VLMs. Overall, there has been increasing attention paid to prompt learning due to its potential to perform significantly better than zero-shot transfer with a few sets of labelled data.
@@
Figure 1: There are possible candidate labels corresponding to the picture of a falcon. In the right portion of the figure, candidate labels are coloured in blue. In the candidate label set, the hidden true label is underlined to be distinct from the false-positive labels.
@@
While prompt learning has demonstrated effectiveness and efficiency in few-shot supervised learning, the true labels must be provided for the training data used in prompt learning. This is a significant defect and will limit the usage of prompt learning in various real-world scenarios because we may be unable to collect accurate labels due to security issues or labelling difficulties. Fortunately, obtaining a set of candidate labels that includes the true label in these situations is easier. For example, as shown in Figure 1, it is challenging to determine which is the true label from ‘Hawk’, ‘Eagle’, and ‘Falcon’, hence all these three labels can be considered as candidate labels during the annotation process of this bird recognition task. As we see, learning with only candidate labels (also widely known as partial-label learning [44, 47, 30, 11]) is practically significant, which also has arisen in many vital applications such as web mining [28], online annotation [41] and ecoinformatics [27]. Nevertheless, existing PLL methods primarily focus on training a model from scratch, and the effectiveness of PLL in the new training paradigm called prompt learning remains unconfirmed. To bridge this research gap, we, for the first time, explore the validity and potential approaches for prompt learning with candidate labels.
This paper provides empirical evidence that prompt learning combined with the prevailing PLL training objectives exhibits superior performance when learning with only candidate labels. Moreover, we conjecture the reason behind the robustness of prompt learning to candidate labels lies in its special position of learnable parameters, which endows the model with strong zero-shot ability, thus mitigating the error accumulation problem [49] in PLL. However, as experimentally suggested, candidate labels will degrade the model’s performance as the level of label ambiguity increases, leading to intolerably poor generalization performance. In previous research, it is reasonable for a PLL method to fail to learn from such highly ambiguous candidate labels (an instance with dozens of labels). However, we believe that the powerful pre-trained VLMs are capable of learning from such ambiguous labels in the procedure of fine-tuning.
@@
Accordingly, for the purpose of leveraging the powerful zero-shot ability of VLMs more properly to guide learning with candidate labels, we propose a simple yet effective framework that dynamically mixes the class posteriors predicted by both the handcrafted and learnable prompt, followed by aligning the mixed class posterior with the model output. Besides, because of the simplicity and flexibility of our framework, it can cooperate with any current PLL training objectives. With our framework, the overall performance of PLL methods has improved by a large margin when fine-tuning VLMs with candidate labels.
@@
  • We provide the first study on the scenario when large vision-language models are fine-tuned with only candidate labels.
  • We demonstrate empirically and explain that prompt learning is more advantageous than other fine-tuning methods, for learning with candidate labels.
  • We propose a framework that guides the process of candidate label disambiguation by aligning the dynamically mixed predicted class posterior of both the handcrafted prompt and the learnable prompt with the model output.
@@
2 Related Work
@@
Vision-language models with fine-tuning. Recently, significant advancements have been made in the field of vision-language models (VLMs) [19, 38, 51, 52]. Unlike models learning from uni-modal supervision, VLMs align visual and textual signals to learn rich representation from massive image-text pairs in the pre-training stage. One of the keys to the excellent performance of VLMs is the tremendous amount of training data. CLIP [38] trains both the text and image encoder simultaneously using a contrastive loss on a dataset of 400 million image-text pairs. ALIGN [19] leverages a noisy pre-training dataset of 1.8 billion pairs to train the model. However, although VLMs have shown promising performance in generalizing to new concepts without auxiliary information, their large amount of parameters makes it impractical to fine-tune the entire model [25]. Besides, full-parameter fine-tuning will also make large models prone to overfitting to downstream tasks and catastrophic forgetting [32]. To address the above issues, multiple transfer learning methods [57, 12, 38, 55] are proposed to adapt VLMs to downstream tasks effectively and efficiently. Linear probe [38] freezes the pre-trained VLM and trains a linear classifier on top of the image encoder. CLIP-adapter [12] introduces an adapter with residual-style feature blending with the pre-trained features. Notably, the success of most of the fine-tuning methods heavily depends on the acquisition of precisely labelled data.
@@
Prompt learning. Among all fine-tuning methods that adapt VLMs to a new task, one typical approach is prompt learning. It treats the text prompt, e.g., “a photo of a <CLS>” as continuous learnable parameters, and optimizes the prompt with multiple vision tasks, including image classification [57, 56, 22, 50, 20], dense prediction [39, 13, 26], etc. CoOp [57] is the first work that migrated prompt learning to vision tasks, which fine-tunes CLIP by optimizing the parameters of the learnable textual prompt (also called soft prompt) while keeping the class token fixed. MaPLe [22] adds multi-modal prompts and learns them mutually to align both representation spaces dynamically. Moreover, recent studies have revealed the advantages of prompt learning when fine-tuning with weakly supervised or unsupervised data [46, 33, 17, 16]. Wu et al. [46] demonstrates that the process of prompt learning is robust to label noise. UPL [17] utilizes handcrafted prompts to generate pseudo labels on unlabelled data and adopts the confident examples per class to tune the learnable prompt.
Partial-label learning. Partial-label learning (PLL) allows each training example to be annotated with a candidate set, containing the true label. Two mainstream strategies have been developed to address this problem: the averaged-based strategy and the identification-based strategy. Average-based strategy disambiguates the candidate labels by treating them equally [18, 4, 54]. Identification-based strategy regards the true label as a latent variable and selects the most likely label for training [21, 9, 34, 10], which is prone to error accumulation if the wrong label is selected initially [49]. As deep learning thrives, many PLL algorithms have been proposed for training with deep neural network [11, 30, 29, 53]. Feng et al. [11] assumes the generation process of partial labels and derives classifier-consistent and risk-consistent methods from a theoretical perspective. PRODEN [30] updates the model parameters and identifies the true labels seamlessly. PiCO [44] divides the learning process into representation learning by contrastive loss and label disambiguation by prototype and pseudo target updating. In particular, as the scale of the parameters and training data of modern deep neural expands [6, 2, 23], the deep learning community has embraced a new training paradigm of pre-training and fine-tuning. Nonetheless, to the best of our knowledge, little research has explored PLL under this new training paradigm.
3 Preliminaries of Prompt Learning
In this work, we show that prompt learning is a preferred choice when tuning VLMs with candidate labels and propose a novel framework to augment its performance. Primarily, we conduct the experiments and build our approach upon a prevailing method in prompt learning: CoOp [57], which tunes the textual prompt based on a well-known VLM called CLIP [38]. In this section, we briefly revisit the details of CLIP and CoOp.
CLIP. CLIP (Contrastive Language-Image Pre-training) is a cutting-edge VLM that learns joint image-text representations through contrastive learning. It consists of both an image encoder and a text encoder. The image encoder encodes visual signals, which can be constructed by ResNet [15] or ViT [7]. The text encoder encodes textual signals, which is constructed by Transformer [43]. By projecting data from both modalities into a shared representation space, CLIP aligns them with a contrastive loss function. The contrastive loss encourages pairs of image-text embeddings to be pushed closer, effectively learning a powerful multi-modal understanding mechanism. With this pre-training approach, CLIP achieves remarkable performance across various visual tasks, particularly zero-shot image classification. In the zero-shot image classification setting, an image $\boldsymbol{x}$ is encoded into a normalized feature $\boldsymbol{f}^{v}$. Likewise, on the text encoder side, a handcrafted prompt of the form “a photo of a <CLS>.” is concatenated with the prior class tokens to generate the textual input. After being encoded by the text encoder, assuming the total number of classes is $C$, a group of normalized text embeddings $\{\boldsymbol{f}^{t}_{j}\}_{j=1}^{C}$ is obtained. Finally, the class posterior is estimated as:
$p_{\mathrm{zs}}(y=i|\boldsymbol{x})=\frac{\mathrm{exp}(\mathrm{sim}(\boldsymbol{f}^{v},\boldsymbol{f}_{i}^{t})/\tau)}{{\textstyle\sum_{j=1}^{C}}\mathrm{exp}(\mathrm{sim}(\boldsymbol{f}^{v},\boldsymbol{f}_{j}^{t})/\tau))},$ (1)
where $\mathrm{sim}(\cdot,\cdot)$ denotes the cosine similarity, and $\tau$ is a temperature factor that controls the range of the output logits.
CoOp. Although CLIP is amazing at zero-shot transfer, its performance is sensitive to differently designed prompts. In order to overcome the inefficiency of the handcrafted prompt designs, CoOp (Context Optimization) [57] replaces the fixed sentence: “a photo of a <CLS>.” with learnable vectors that have the same dimension as the word embedding of CLIP. These learnable vectors will be optimized with a few labelled examples. In the remaining sections, we will study CoOp as a typical case of prompt learning for VLMs.
@@ 3 Preliminaries of Prompt Learning
Ultimately, the learnable context vectors $\{\boldsymbol{v}_{k}\}_{k=1}^{M}$ are optimized on a dataset $D=\{(\boldsymbol{x}_{i},y_{i})_{i=1}^{N}\}$ with the cross-entropy loss:
@@
Notably, optimizing this training objective of CoOp requires examples of true labels. But in many realistic scenarios, the true label is not accessible due to multiple reasons. Instead, we can obtain a candidate label set $Y_{i}$ that contains the true label $y_{i}$, i.e., $y_{i}\in Y_{i}$. Unfortunately, we cannot utilize the candidate label set to optimize the above training objective. Therefore, the next section will study prompt learning with candidate labels.
@@ 4 Prompt Learning with Candidate Labels
Figure 2: Illustration of our framework. Our framework includes a prompt alignment module and a PLL method module. Regarding the prompt alignment module, two different prompts are input to the text encoder, including the handcrafted prompt and the learnable prompt. Afterwards, we mix the recalculated class posteriors yielded by the handcrafted prompt and the soft prompt. Then, the mixed class posterior is aligned with the output of the soft prompt using the re-weighted cross-entropy loss. For the PLL method module, any off-the-shelf PLL training objectives can be combined with the prompt alignment module. It is important to know that during fine-tuning, all the parameters of this framework are frozen except for the learnable prompt.
Prompt learning has been shown to be effective when fine-tuning VLMs, provided a few examples with perfect labels. Nevertheless, it remains unclear how prompt learning performs with candidate labels.
In this section, we first empirically demonstrate the superiority of prompt learning with PLL training objectives. Then, we explain why prompt learning performs better with candidate labels. Finally, based on the observations above, we propose a framework that dynamically aligns the mixed class posterior with the model’s output to enhance the performance. Figure 2 depicts our framework.
@@ 4.1 Pilot Experiments
This part explores prompt learning with PLL training objectives and presents our key findings. The most important observation is that prompt learning significantly outperforms linear probe with ambiguous supervision. Linear probe [42, 38] is a strong baseline in few-shot learning, which trains a linear classifier on top of the pre-trained frozen image encoder of VLMs. In the following experiments, we will use linear probe and prompt learning to fine-tune the pre-trained model with different types of PLL training objectives to learn from the candidate labels.
In the pilot experiments, we utilize ResNet-50 [15] as the same visual backbone of the two fine-tuning methods and compare their performance on the datasets of UCF101 and Caltech101 under different levels of label ambiguity. Specifically, We define the level of label ambiguity $q$ as the probability of flipping negative labels $\bar{y}_{i}\neq y_{i}$ to false-positive labels inside the candidate label set $Y_{i}$: $q=\mathrm{Pr}(\bar{y}_{i}\in Y_{i}|\bar{y}_{i}\neq y_{i})$. It is noteworthy that the true label is guaranteed to be inside the candidate label set. For linear probe, the original learning rate is selected from $\{50,25,10,5,2.5,1,0.5,0.25,0.1,0.05,0.0025,0.001\}$ to ensure optimal performance. Moreover, the other training settings of linear probe and prompt learning are consistent with the implementation detail in the experiment section. The result is shown in Figure 3.
@@
Figure 3: Performance comparison with multiple fine-tuning approaches on UCF101 [40] and Caltech101 [8] with candidate labels of the incremental label ambiguity. Prompt learning and linear probe are incorporated with the PLL training objective of PiCO [44]. vPLL means a simple baseline that treats every candidate label as the ground-truth label and uses cross-entropy loss to learn.
As seen from Figure 3, although linear probe and prompt learning both achieve excellent performance with true labels, they suffer from performance drop with candidate labels. However, prompt learning is apparently more robust and performs better under all circumstances. In addition, we add a vanilla method for comparison: ‘vPLL’, which treats every candidate label as the ground-truth label and uses cross-entropy loss to learn from them. Compared with prompt learning + PLL, prompt learning + vPLL performs much worse, showing that the PLL training objective has played a significant part in enabling VLMs to learn effectively from candidate labels.
To summarize, this result shows that prompt learning with PLL training objectives can perform well with candidate labels. However, if the dataset is partialized with a higher level of label ambiguity, the performance of prompt learning will get intolerably worse. (-4.4% for Caltech101 and -11.3% for UCF101)
The following part will be used to explain why prompt learning is more robust with ambiguous labels than linear probe.
4.2 Reasons for the Robustness of Prompt Learning
In order to investigate the reasons behind the excellent performance of prompt learning, the test accuracy of three models is evaluated by epoch on datasets of UCF101 and Caltech101 when $q=0.3$, as shown in Figure 4. The models are:
  • Prompt learning with random initialization.
  • Prompt learning with a handcrafted initialization (i.e., handcrafted prompt).
  • Linear probe
To be clear, the handcrafted prompts of the second variant are chosen from the result of the prompt engineering from CLIP [38].
Figure 4: We record the test accuracy of each epoch of different fine-tuning methods with the PLL training objective: PiCO [44]. Prompt Learning* refers to prompt learning, whose soft prompt is initialized with a handcrafted prompt.
From Figure 4, we can see that, for prompt learning, the label ambiguity will not sabotage the generalization ability that the VLM initially possesses. Moreover, with a proper PLL training objective, the model can quickly adapt to these training data. Furthermore, even the performance of the model with a randomly initialized prompt can rapidly return to zero-shot performance. Conversely, due to the lack of zero-shot ability, linear probe struggles with the candidate labels at first and cannot learn well from the candidate labels.
In contrast to linear probe, the special position of the tuning parameter of prompt learning keeps the zero-shot ability of VLMs. Even if the tunable prompt is initialized randomly, due to the regularization effect of the fixed class token, the model is more robust to the noisy false-positive labels [46] and will achieve the zero-shot ability to the downstream data through only a few epochs. Remarkably, the zero-shot ability is essential in candidate label disambiguation because it enables the model to select a more likely true label initially, which will mitigate the error accumulation problem [49] in PLL. The error accumulation problem means that since most PLL methods leverage the model’s prediction from the previous iteration, the bias will be accumulated and pose a challenge for adjusting mistakes once a false-positive label is identified as the true label, thus severely hurting the generalization of the model. For linear probe, it can only randomly guess which label is correct at first and will significantly suffer from the error accumulation problem. As a result, unlike prompt learning, the pre-trained model’s powerful representation ability does not effectively contribute to label disambiguation and will later be corrupted by them.
However, we notice that as training proceeds or the label ambiguity increases, the model performance will inevitably be degraded by the candidate labels in prompt learning. Hence, we propose a framework that explicitly utilizes the zero-shot ability of VLMs to learn better from candidate labels.
4.3 The Proposed Framework

### v3

@@ Abstract
Vision-language models (VLMs) can learn high-quality representations from a large-scale training dataset of image-text pairs. Prompt learning is a popular approach to fine-tuning VLM to adapt them to downstream tasks. Despite the satisfying performance, a major limitation of prompt learning is the demand for labeled data. In real-world scenarios, we may only obtain candidate labels (where the true label is included) instead of the true labels due to data privacy or sensitivity issues. In this paper, we provide the first study on prompt learning with candidate labels for VLMs. We empirically demonstrate that VLMs can learn from candidate labels through prompt learning. Nonetheless, its performance drops when the label ambiguity increases. In order to improve its robustness, we propose a simple yet effective framework that better leverages the prior knowledge of VLMs to guide the learning process with candidate labels. Specifically, our framework disambiguates candidate labels by aligning the model output with the mixed class posterior jointly predicted by both the learnable and the handcrafted prompt. Besides, our framework can be equipped with various training objectives for learning with candidate labels to further improve their performance. Extensive experiments demonstrate the effectiveness of our proposed framework.
@@
Multimodal Models Prompt Learning Candidate Labels.
@@ 1 Introduction
Vision-language models (VLMs) such as CLIP [32], ALIGN [16], and Coca [42] have become excellent base models in multiple domains, most of which employ a dual-encoder architecture to align the natural images with descriptive texts. Remarkably, this special training pattern has endowed VLMs with superior zero-shot transfer performance on visual recognition tasks. In specific, during the inference, the pre-trained text encoder receives inputs in the form of man-crafted prompts, e.g., “a photo of <CLS>.”. Subsequently, all the generated textual embeddings are matched with the visual embedding obtained from the image encoder to predict the image category. However, the powerful zero-shot ability of VLMs was shown to be heavily dependent on the choice of handcrafted prompts, which needs substantial efforts and professional domain knowledge to design [49]. To avoid the manual design of the prompts, prompt learning [49] is proposed, which treats the textual prompt as additional learnable parameters and tunes them while keeping all the original parameters of the pre-trained model fixed. Later, the concept of the prompt is extended to visual prompt [17] and multi-modal prompt [19] in VLMs. Overall, there has been increasing attention paid to prompt learning due to its potential to perform significantly better than zero-shot transfer with a few sets of labeled data [49, 48, 50, 19, 17].
@@
Figure 1: There are possible candidate labels corresponding to the image of a falcon. In the right portion of the figure, candidate labels are colored in blue. In the candidate label set, the hidden true label is underlined to be distinct from the false-positive labels.
@@
While prompt learning has demonstrated effectiveness and efficiency in few-shot supervised learning, the true labels must be provided for the training data used in prompt learning. This is a significant defect and will limit the usage of prompt learning in various real-world scenarios because we may be unable to collect accurate labels due to security issues or labeling difficulties. Fortunately, obtaining a set of candidate labels that includes the true label in these situations is easier. For example, as shown in Figure 1, it is challenging to determine which is the true label from ‘Hawk’, ‘Eagle’, and ‘Falcon’, hence all these three labels can be considered as candidate labels during the annotation process of this bird recognition task. As we see, learning with only candidate labels (also widely known as partial-label learning (PLL) [37, 40, 26, 10]) is practically significant, which also has arisen in many vital applications such as web mining [24], online annotation [35] and ecoinformatics [23]. Nevertheless, existing PLL methods primarily focus on training a model from scratch, and the effectiveness of PLL in the new training paradigm called prompt learning remains unconfirmed. To bridge this research gap, we, for the first time, explore the validity and potential approaches for prompt learning with candidate labels.
This paper empirically shows that prompt learning combined with the prevailing PLL training objectives in a vanilla way can learn from candidate labels. However, as experimentally suggested, if the candidate labels become more ambiguous, the model’s performance will drop significantly. Fortunately, prompt learning is still more robust than linear probe [32], another tuning method that trains a linear classifier on top of a frozen pre-trained model. We conjecture that the reason lies in the prior knowledge brought by the fixed class token, which can keep VLMs from over-fitting to the false-positive labels in the candidate label set and provide VLMs with preferred zero-shot ability, thus mitigating the error accumulation problem [41] in PLL when the label ambiguity increases.
@@
Therefore, to enhance the robustness of prompt learning with candidate labels, we propose a simple yet effective framework incorporating the handcrafted prompt to distill the model with more comprehensive prior knowledge. Concretely, it dynamically mixes the class posteriors predicted by both the handcrafted and learnable prompt, followed by aligning the mixed class posterior with the model output. Besides, due to the simplicity and flexibility of our framework, it can cooperate with any current PLL training objectives. With our framework, the overall performance of various PLL training objectives has improved by a large margin when tuning VLMs with candidate labels.
@@
  • We provide the first study on the scenario when vision-language models are tuned with only candidate labels.
  • We demonstrate empirically and explain that prompt learning combined with PLL training objectives in a vanilla way can learn from candidate labels but is not robust when the label ambiguity is high.
  • A framework is proposed to enhance the robustness of prompt learning with candidate labels by aligning the model output with the dynamically mixed prediction by both the handcrafted and the learnable prompt.
@@
Figure 2: Illustration of our framework. Our framework includes a prompt alignment module and a PLL method module. Regarding the prompt alignment module, two different prompts are input to the text encoder, including the handcrafted prompt and the learnable prompt. Afterwards, we mix the recalculated class posteriors yielded by the handcrafted prompt and the learnable prompt. Then, the mixed class posterior is aligned with the output of the learnable prompt using the re-weighted cross-entropy loss. For the PLL method module, any prevailing PLL methods can be combined with the prompt alignment module. It is important to know that during fine-tuning, all the parameters of this framework are frozen except for the learnable prompt.
2 Preliminaries of Prompt Learning and PLL
2.1 Prompt Learning
@@
Recently, significant advancements have been made in the field of vision-language models (VLMs) [16, 32, 42, 43]. Unlike models learning from uni-modal supervision, VLMs align visual and textual signals to learn rich representation from massive image-text pairs in the pre-training stage. One of the keys to the excellent performance of VLMs is the tremendous amount of training data. CLIP [32] trains both the text and image encoder simultaneously using a contrastive loss on a dataset of 400 million image-text pairs. ALIGN [16] leverages a noisy pre-training dataset of 1.8 billion pairs to train the model. However, although VLMs have shown promising performance in generalizing to new concepts without auxiliary information, their large amount of parameters makes it impractical to fine-tune the entire model [22]. Besides, full-parameter fine-tuning will also make large models prone to overfitting to downstream tasks and catastrophic forgetting [28]. To address the above issues, multiple transfer learning methods [49, 11, 32, 47] are proposed to adapt VLMs to downstream tasks effectively and efficiently. Linear probe [32] freezes the pre-trained VLM and trains a linear classifier on top of the image encoder. CLIP-adapter [11] introduces an adapter with residual-style feature blending with the pre-trained features. Notably, most fine-tuning methods heavily depend on the acquisition of precisely labeled data.
@@
Among all fine-tuning methods that adapt VLMs to a new task, one typical approach is prompt learning. It treats the text prompt, e.g., “a photo of a <CLS>” as continuous learnable parameters, and optimizes the prompt with multiple vision tasks, including image classification [49, 19], dense prediction [33], ‘. CoOp [49] is the first work that migrated prompt learning to vision tasks, which tunes CLIP by optimizing the parameters of the learnable textual prompt (also called soft prompt) while keeping the class token fixed.
@@
At last, the learnable context vectors $\{\boldsymbol{v}_{k}\}_{k=1}^{M}$ are optimized on a dataset $D=\{(\boldsymbol{x}_{i},y_{i})_{i=1}^{N}\}$ with the cross-entropy loss:
@@
Notably, optimizing this training objective of CoOp requires examples of true labels. But in many realistic scenarios, the true label is not accessible due to multiple reasons. The former research has focused on prompt learning with noisy labels [39] or no label [14, 45] and showed prompt learning is not only robust with label noise but also can effectively learn from unlabelled data with specifically designed algorithms.
In this work, we focus on the scenario where only a candidate label set can be obtained, which we unfortunately cannot utilize to optimize the above training objective. Therefore, we will study prompt learning with candidate labels using PLL training objectives instead of cross-entropy loss.
2.2 Partial-label Learning
Partial-label learning (PLL) allows each training example to be annotated with a candidate set, containing the true label. The problem is defined as follows: In the training process, we can obtain a candidate label set $Y_{i}$ for sample $x_{i}$ that contains the true label $y_{i}$, i.e., $y_{i}\in Y_{i}$. To simplify the problem, in the following experiments, unless otherwise stated, we assumed that the the candidate labels in partial label datasets manually constructed (also called ‘partialized’) from fully-supervised datasets for PLL method evaluation are generated uniformly by flipping negative labels $\bar{y}_{i}\neq y_{i}$ to false-positive labels inside the candidate label set with label ambiguity $q=P(\bar{y}_{i}\in Y_{i}|\bar{y}_{i}\neq y_{i})$. Two mainstream strategies have been developed to address this problem: the averaged-based strategy and the identification-based strategy. Average-based strategy disambiguates the candidate labels by treating them equally [15, 4, 46]. However, the main drawback of this strategy is that the learner can be seriously affected by multiple false-positive labels [25]. Identification-based strategy regards the true label as a latent variable and selects the most likely label for training [18, 8, 29, 9], which is prone to error accumulation if the wrong label is selected initially [41]. As deep learning thrives, many PLL algorithms have been proposed for training with deep neural network [10, 26, 25, 44]. Feng et al. [10] assumes the generation process of partial labels and derives classifier-consistent and risk-consistent methods from a theoretical perspective. PRODEN [26] updates the model parameters and identifies the true labels seamlessly. PiCO [37] divides the learning process into representation learning by contrastive loss and label disambiguation by prototype and pseudo-target updating. In particular, as the scale of the parameters and training data of modern deep neural expands [6, 2, 20], the deep learning community has embraced a new training paradigm of pre-training and fine-tuning. Recently, Zhang et al. [45] proposed a strategy that allows VLMs to generate candidate pseudolabels to enhance prompt learning with unlabelled data. Their work is orthogonal to ours, as we enhance prompt learning performance with human-annotated candidate labels, while they focus on using unlabelled data.
@@ 3 Prompt Learning with Candidate Labels
Prompt learning has been shown to be effective when tuning VLMs, provided a few examples with perfect labels. Nevertheless, it remains unclear how prompt learning performs with candidate labels.
In this section, we empirically demonstrate the viability of it combined with PLL training objectives in a vanilla way for learning with candidate labels and explain the reasons. Then, based on the observations above, we propose a framework that dynamically aligns the mixed class posterior with the model’s output to enhance performance. Figure 2 depicts our framework.
@@ 3.1 Pilot Experiments
Figure 3: Performance comparison with multiple fine-tuning approaches combined with PiCO [37] in a vanilla way on UCF101 [34] and Caltech101 [7] with candidate labels of the incremental label ambiguity. vPLL means a simple baseline that treats every candidate label as the ground-truth label and uses cross-entropy loss to learn.
This part explores prompt learning with PLL training objectives and presents our key findings. Linear probe [36, 32] is a strong baseline in few-shot learning, which trains a linear classifier on top of the pre-trained frozen image encoder of VLMs. PiCO [37], a prevailing PLL training objective, uses embedding prototypes to identify the ground-truth label from the candidate set and train the classifier, which in turn selects a positive set to calculate the contrastive loss and better align the embedding prototypes.
In the pilot experiment, we use linear probe and prompt learning to tune the pre-trained model simply combined with PiCO on the uniformly partialized datasets of UCF101 [34], and Caltech101 [7].
Moreover, we utilize ResNet-50 [13] as the same visual backbone of the two fine-tuning methods and compare their performance on the datasets of UCF101 and Caltech101 under different levels of label ambiguity. For linear probe, the learning rate is selected from {2.5,1,0.5,0.25,0.1,0.05, 0.01} to ensure optimal performance. The other training settings for linear probe and prompt learning are consistent with the implementation details in the experiment section. Additionally, in the remaining content, ‘prompt learning combined with PLL training objectives in a vanilla way’ is abbreviated as ‘vanilla prompt learning’.
The result is shown in Figure 3. We can see that vanilla prompt learning outperforms CLIP zero-shot, demonstrating the viability of learning with candidate labels through prompt learning. However, as the level of label ambiguity increases from 00 to $0.5$, the performance deteriorates significantly (-4.4% for Caltech101 and -11.3% for UCF101), which is unacceptable.
On the other hand, it is shown that vanilla prompt learning is more robust with candidate labels than linear probe. We conjecture that it is the fixed class token that introduces the prior knowledge, enhancing the model’s robustness. Specifically, the prior knowledge not only acts as a regularizer, keeping the model from over-fitting to the false-positive labels [39], but also equips the model with a preferred zero-shot ability to identify the concealing true label better. These effects can jointly mitigate the error accumulation problem [41] in PLL.
Therefore, to enhance the model’s robustness with candidate labels, inspired by the advantages of the fixed class token, we propose a framework that introduces the handcrafted prompt to distill the model with full prior knowledge.
3.2 Our Framework